# sources.bib

### v1

@@ 4 The Chirality Decomposition
By using the soldering form, the nature of this decomposition is algebraic and occurs if and only if the underlying spin representation is of complex type. In this case, the structure group of $\mathcal{F}^{\mathbb{C}}$ is $\operatorname{Spin}^{\mathbb{C}}(V,q)$ and embeds canonically into $\operatorname{GL}_{\mathbb{C}}(S^{\mathbb{C}})$. It follows that the almost complex structure integrates to a complex structure along $\mathcal{F}^{\mathbb{C}}$. This means that $\mathcal{F}^{1,0}$ and $\mathcal{F}^{0,1}$ can be interpreted as holomorphic and antiholomorphic distributions, respectively.
@@
Due to the non-integrability conditions on $\mathcal{F}$, we do not try to further decompose the exterior differential on forms into holomorphic and antiholomorphic parts. Rather, we consider a spinorial connection $\varphi^{s}\in\mathcal{A}_{\hat{Q}}$, which preserves the bosonic resp. spinorial distributions (note that the spinorial connection is inherently defined on the principal bundle $\hat{Q}\to M$ of $\operatorname{Spin}^{\mathbb{C}}(V,q)$ frames of the complexified tangent sheaf $\mathcal{T}_{M}^{\mathbb{C}}$). The corresponding covariant exterior derivative decomposes globally as
@@
as the $\mathbb{C}$-structure of $\mathcal{F}^{\mathbb{C}}$ commutes with the spinorial structure. Here, the suggestive latter notation is chosen due to the analogy with the Dolbeault operators on complex manifolds
@@
It is clear that both $M^{1|0}$ and $M^{0|1}$ constitute locally ringed spaces. We prove now that $M^{1|0}$ is a CS manifold (the case of $M^{0|1}$ is analogous). By lemma 17, it suffices to check that the distribution $\mathcal{B}^{\mathbb{C}}\oplus\mathcal{F}^{1,0}$ is integrable. By Frobenius’ theorem, this amounts to verifying that the Frobenius pairing vanishes, that is, the $\mathcal{F}^{0,1}$ components of the commutator are zero. By means of the soldering form, the Frobenius obstructions of the kind $\operatorname{pr}_{0|1}([\cdot,\cdot])$ can be lifted to the bundle of frames, where they take the form of symmetric morphisms of $\operatorname{Spin}(V,q)$ representations $S^{1,0}\otimes S^{1,0}\to S^{0,1}$, $V^{\mathbb{C}}\otimes V^{\mathbb{C}}\to S^{0,1}$, and $V^{\mathbb{C}}\otimes S^{1,0}\to S^{0,1}$. However, these morphisms are all identically zero:
@@
Local freeness is clear by local freeness vector bundles and because the complex structure gives non-degenerate algebraic relations on the sections that smoothly depend on $M$. To prove the first claim, let $X\in\Gamma^{\omega}(M,E)$ be a chiral section and let $f\in\mathcal{O}^{\omega}_{M}$ be a chiral function. We compute
@@
If the spinorial representation $S^{1,0}$ underlying $\mathcal{F}^{1,0}$ carries an invariant non-degenerate skew-symmetric bilinear form, the CS manifold $M^{1,0}$ carries a canonical volume form, denoted $\mathrm{vol}^{1,0}\in{\mathcal{B}}er_{M}^{1,0}$. The analogous statement holds for $M^{0,1}$.
@@ 5 Restriction to Ordinary Spacetimes
$i^{*}Q^{\mathcal{B}}\times_{\check{M}}i^{*}Q^{\mathcal{F}}\to\check{M}.$ (45)
The first factor defines the Riemannian structure, and the latter one defines the $\operatorname{Spin}^{\mathbb{F}}(V,q)$ structure. By definition of the superspace structure, there exists a map of principal bundles $Q^{\mathcal{F}}\to Q^{\mathcal{B}}$ with kernel $K$, and pulling back by $i^{*}$, we obtain the compatible structures
$i^{*}Q^{\mathcal{B}}\to i^{*}Q^{\mathcal{F}}\to\check{M}.$ (46)
Note that the pullback $i^{*}Q^{\mathcal{F}}$, together with the spinorial representation $S$, allows for the construction of a spinor bundle over the ordinary spacetime
$S\check{M}=i^{*}Q^{\mathcal{F}}\times_{\operatorname{Spin}^{\mathbb{F}}(V,q)}\Pi S\to\check{M}.$ (47)
@@
Finally, we have the following relations between the spinorial connection on $M$ and the spin connection on $i^{*}Q^{\mathcal{F}}\to\check{M}$:
@@
Let $\varphi^{s}\in\mathcal{A}^{s}_{Q}$ be a spinorial connection. The pullback connection $i^{*}\varphi^{s}\in\mathcal{A}_{i^{*}Q}$ is a $\operatorname{Spin}^{\mathbb{F}}(V,q)$-connection on $i^{*}Q=\check{Q}$ (and thus in particular on the spinor bundle $S\check{M}$); and the pullback $i^{*}\left((\varphi^{s})^{\mathcal{B}}\right)\in\mathcal{A}_{i^{*}Q^{\mathcal{B}}}$ of the $Q^{\mathcal{B}}$-part of $\varphi^{s}$ is the Levi-Civita connection on $(\check{M},g|_{\check{M}})$.
@@ 6.1 The Construction
[…]
@@
$\delta_{\mathcal{F}}:\tau^{*}\Pi S\check{M}\to T\Pi S\check{M},\quad X\mapsto X+\frac{1}{2}\nabla^{\phi}_{\Gamma(\mathcal{E},X)},$ (68)
@@
as upon restricting to the body, the structure sheaf is reduced to ordinary functions on $\check{M}$, where the covariant derivative reduces to the ordinary differential. The term involving $[\nabla_{\Gamma(\mathcal{E},\partial_{\mu})}^{\varphi},\nabla_{\Gamma(\mathcal{E},\partial_{\nu})}^{\varphi}]$ vanishes along the body as it has coefficients that are linear in $\Pi S$. Putting everything together, we find
@@
Let us fix (global) generators $\{\partial_{i}\}$ corresponding to a basis $\{e_{i}\}$ of $V$, and $\{\partial_{\mu}\}$ corresponding to a basis $s_{\mu}$ of $\Pi S$. Let ${\Gamma_{\mu\nu}^{i}}v_{i}=\Gamma(s_{\mu},s_{\nu})$. In these coordinates, the structure distribution is given by
@@
such that $[D_{\mu},D_{\nu}]={\Gamma_{\mu\nu}}^{i}\partial_{i}$. This affine superspace constitutes a local model for more general spinorial superspaces.
@@ 6.2 Computing Integrals
where the $(e_{i})$ are a frame on $\check{M}$, i.e., a section of $\check{Q}\to\check{M}$. In what follows, we identify frames and coframes by dualising the bases of the respective distributions. Note that there exists a $\operatorname{GL}(m|n)$ transformation of unit Berezinian relating $q$ with a frame that has the $\partial_{\mu}$ as odd generators (rather than $\delta_{\mathcal{F}}(\partial_{\mu}$), given explicitly by
@@
$\int_{M}f[q]=\int_{\check{M}}i^{*}(\partial_{1}\dots\partial_{n}f)e_{1}\wedge\dots\wedge e_{m}$ (88)
@@ Appendix B Frame Computations
Let $\psi$ be the sYM spinor associated to a connection $\mathsf{a}\in\mathcal{A}_{P}$, let $\varphi^{s}\in\mathcal{A}_{\hat{Q}}^{s}$ be a spinorial connection, and let $X$ be the horizontal lift of a section of the spinorial distribution $\mathcal{F}$ to $\hat{Q}\times_{M}P$. Denote by $A=\mathsf{a}\oplus\varphi^{s}$ the product connection, set $\nabla^{A}\eqqcolon\nabla$ and let $i:\check{M}\to M$ be the embedding of the ordinary spacetime manifold. It holds
@@
$\sqrt{n}\nabla_{X}\psi$ $=\iota_{X}\mathrm{d}^{A}\langle\phi,\tilde{\gamma}\rangle=\langle\iota_{X}\mathrm{d}^{A}\phi,\tilde{\gamma}\rangle,$ (126)
@@
$\sqrt{n}\nabla_{X}\psi$ $=\left\langle\iota_{X}\mathrm{d}^{A}\phi,\sum_{Y}\tilde{\gamma}(Y^{\vee})\otimes Y\right\rangle$ (128)
$=\sum_{Y}\langle\iota_{Y}\iota_{X}\mathrm{d}^{A}\phi,\tilde{\gamma}(Y^{\vee})\rangle$ (129)
@@
$=i^{*}\langle\mathrm{d}^{a}\phi,\tilde{\gamma}\circ\gamma\circ\tilde{\epsilon}\rangle,$ (143)
@@
Let $\Delta_{\epsilon}^{A}=\epsilon^{XY}\nabla_{X}^{\circ}\nabla_{Y}$, let $a=i^{*}\mathsf{a}$, and let $\check{\psi}=i^{*}\psi$. It holds
$\big{(}\Delta^{A}_{\epsilon}\psi\big{)}\big{|}_{\check{M}}=2D\!\!\!\!/\,^{a}\lambda.$ (144)
@@
$\sqrt{n}\cdot\Delta^{A}_{\epsilon}\psi$ $=\epsilon^{XY}\iota_{X}\mathrm{d}^{A}\iota_{Y}\mathrm{d}^{A}\langle\phi,\tilde{\gamma}\rangle$ (145)
$=\epsilon^{XY}\langle\iota_{Z}\iota_{X}\mathrm{d}^{A}\iota_{Y}\mathrm{d}^{A}\phi,\tilde{\gamma}(Z^{\vee})\rangle$ (146)
$=\epsilon^{XY}\langle\iota_{X}\iota_{Z}L_{Y}\mathrm{d}^{A}\phi,\tilde{\gamma}(Z^{\vee}),\rangle$ (147)
$=\epsilon^{XY}\langle\iota_{X}(\iota_{[Z,Y]}-\mathrm{d}^{A}\iota_{Y}\iota_{Z})\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}(Z^{\vee})\rangle$ (148)
$=-\epsilon^{XY}\langle\iota_{[Z,Y]}\iota_{X}\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}(Z^{\vee})\rangle-\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{A}\iota_{Y}\iota_{Z}\mathrm{d}^{A}\phi,\tilde{\gamma}(Z^{\vee})\rangle$ (149)
@@
$\epsilon^{XY}\langle\iota_{[Z,Y]}\iota_{X}\mathrm{d}^{A}\phi,\tilde{\gamma}(Z^{\vee})\rangle\big{|}_{\check{M}}$ $=\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{A}\phi,\tilde{\gamma}(Z^{\vee})\otimes\Gamma(Z,Y)\rangle\big{|}_{\check{M}}$ (150)
$=\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{A}\phi,\tilde{\gamma}\circ\gamma(Y)\rangle\big{|}_{\check{M}}$ (151)
$=-\langle\mathrm{d}^{A}\phi,\tilde{\gamma}\circ\gamma\circ\tilde{\epsilon}\rangle\big{|}_{\check{M}}$ (152)
@@
$-\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{A}\iota_{Y}\iota_{Z}\mathrm{d}^{A}\phi,\tilde{\gamma}(Z^{\vee})\rangle\big{|}_{\check{M}}$ $=\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{A}\iota_{[Y,Z]}f,\tilde{\gamma}(Z^{\vee})\rangle\big{|}_{\check{M}}$ (154)
@@
$\Delta^{A}_{\epsilon}\psi\big{|}_{\check{M}}=2D\!\!\!\!/\,^{a}\lambda,$ (158)
@@
$\sqrt{n}\nabla_{X}\chi$ $=\iota_{X}\mathrm{d}^{A}\langle\zeta,\tilde{\gamma}\rangle$ (162)
$=\sum_{Y}\langle\iota_{Y}\iota_{X}\mathrm{d}^{A}\zeta,\tilde{\gamma}(Y^{\vee})\rangle$ (163)
$=-\sum_{Y}\langle\iota_{Y}\iota_{X}\mathrm{d}^{A}(f+\bar{\zeta}),\tilde{\gamma}(Y^{\vee})\rangle$ (164)
$=-\sum_{Y}\langle\iota_{[Y,X]}\mathrm{d}^{A}(f+\bar{\zeta}),\tilde{\gamma}(Y^{\vee})\rangle$ (165)
@@
$\sqrt{n}\nabla_{X}\chi=\iota_{X}\mathrm{d}^{A}\langle\zeta,\tilde{\gamma}\rangle|_{\check{M}}=-\langle f+\bar{\zeta},\tilde{\gamma}\circ\gamma(X)\rangle|_{\check{M}}=-\langle f,\tilde{\gamma}\circ\gamma(X)\rangle|_{\check{M}}$ (166)
@@
$\big{(}\Delta^{A}_{\epsilon}\chi\big{)}\big{|}_{\check{M}}\coloneqq\Big{(}\epsilon^{XY}\nabla_{X}\nabla_{Y}\chi\Big{)}\Big{|}_{\check{M}}=-D\!\!\!\!/\,^{a}\big{(}\chi|_{\check{M}}\big{)}.$ (168)
@@
$\sqrt{n}\cdot\Delta^{A}_{\epsilon}\chi$ $=\epsilon^{XY}\iota_{X}\mathrm{d}^{A}\iota_{Y}\mathrm{d}^{A}\langle\zeta,\tilde{\gamma}\rangle$ (169)
$=\epsilon^{XY}\langle\iota_{Z}\iota_{X}\mathrm{d}^{A}\iota_{Y}\mathrm{d}^{A}\zeta,\tilde{\gamma}(Z^{\vee})\rangle$ (170)
$=\epsilon^{XY}\langle\iota_{X}\iota_{Z}L_{Y}\mathrm{d}^{A}\zeta,\tilde{\gamma}(Z^{\vee}),\rangle$ (171)
$=\epsilon^{XY}\langle\iota_{X}(\iota_{[Z,Y]}-\mathrm{d}^{A}\iota_{Y}\iota_{Z})\mathrm{d}^{A}\zeta,\tilde{\gamma}(Z^{\vee})\rangle$ (172)
$=-\epsilon^{XY}\langle\iota_{[Z,Y]}\iota_{X}\mathrm{d}^{A}\zeta,\tilde{\gamma}(Z^{\vee})\rangle-\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{A}\iota_{Y}\iota_{Z}\mathrm{d}^{A}\zeta,\tilde{\gamma}(Z^{\vee})\rangle$ (173)
@@
$\epsilon^{XY}\langle\iota_{[Z,Y]}\iota_{X}\mathrm{d}^{A}\zeta,\tilde{\gamma}(Z^{\vee})\rangle\big{|}_{\check{M}}$ $=\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{A}\zeta,\tilde{\gamma}(Z^{\vee})\otimes\Gamma(Z,Y)\rangle\big{|}_{\check{M}}$ (174)
$=\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{A}\zeta,\tilde{\gamma}\circ\gamma(Y)\rangle\big{|}_{\check{M}}$ (175)
$=-\langle\mathrm{d}^{A}\zeta,\tilde{\gamma}\circ\gamma\circ\tilde{\epsilon}\rangle\big{|}_{\check{M}}$ (176)
@@
$-\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{A}\iota_{Y}\iota_{Z}\mathrm{d}^{A}\zeta,\tilde{\gamma}(Z^{\vee})\rangle\big{|}_{\check{M}}$ $=\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{A}\iota_{[Y,Z]}(f+\bar{\zeta}),\tilde{\gamma}(Z^{\vee})\rangle\big{|}_{\check{M}}$ (178)

### v2

@@ 4 The Chirality Decomposition
By using the approach via the bundle of frames, the nature of this decomposition can be expressed algebraically and occurs if and only if the underlying spin representation is of complex type. In this case, the structure group of $\mathcal{F}^{\mathbb{C}}$ is $\operatorname{Spin}^{\mathbb{C}}(V,q)$ and embeds canonically into $\operatorname{GL}_{\mathbb{C}}(S^{\mathbb{C}})$. Since the complex nature of the CS manifold $(M,\mathcal{O}_{M}\otimes\mathbb{C})$ only enters along the odd directions (and is therefore to be regarded as a completely algebraic property of the structure sheaf), the almost complex structure can be viewed as a complex structure along $\mathcal{F}^{\mathbb{C}}$. This means that $\mathcal{F}^{1,0}$ and $\mathcal{F}^{0,1}$ can be interpreted as holomorphic and antiholomorphic distributions, respectively.
@@
Due to the non-integrability conditions on $\mathcal{F}$, we do not try to further decompose the exterior differential on forms into holomorphic and antiholomorphic parts. Rather, we consider a spinorial connection $\varphi^{s}\in\mathcal{A}_{\hat{Q}}$. Since $\varphi^{s}$ is a $\operatorname{Spin}^{\mathbb{C}}(V,q)$ connection, and $S^{\mathbb{C}}=S^{1,0}\oplus S^{0,1}$ as $\operatorname{Spin}^{\mathbb{C}}(V,q)$ representations, such a spinorial connection preserves the bosonic resp. spinorial distributions. The corresponding covariant exterior derivative thus decomposes globally as
@@
on fibre bundles over $M$ arising from the decomposition $\mathcal{T}_{M}^{\mathbb{C}}=\mathcal{B}^{\mathbb{C}}\oplus\mathcal{F}^{0,1}\oplus\mathcal{F}^{0,1}$. By this notation, we mean that
$\partial^{\varphi^{s}}:\Omega^{\bullet}(M,F^{1,0})\to\Omega^{\bullet+1}(M,F^{1,0}),$ (57)
and likewise for the other components. The suggestive notation is chosen due to the analogy with the Dolbeault operators on complex manifolds.
@@
It is clear that both $M^{1|0}$ and $M^{0|1}$ constitute locally ringed spaces. We prove now that $M^{1|0}$ is a CS manifold (the case of $M^{0|1}$ is analogous). By lemma 17, it suffices to check that the distribution $\mathcal{B}^{\mathbb{C}}\oplus\mathcal{F}^{1,0}$ is integrable. By Frobenius’ theorem, this amounts to verifying that the Frobenius pairing vanishes, that is, the $\mathcal{F}^{0,1}$ components of the commutator are zero. As in the discussion of Equation (26), the Frobenius obstructions of the kind $\operatorname{pr}_{0|1}([-,-])$ can be lifted to the bundle of frames, where they take the form of symmetric morphisms of $\operatorname{Spin}(V,q)$ representations $S^{1,0}\otimes S^{1,0}\to S^{0,1}$, $V^{\mathbb{C}}\otimes V^{\mathbb{C}}\to S^{0,1}$, and $V^{\mathbb{C}}\otimes S^{1,0}\to S^{0,1}$. However, these morphisms are all identically zero:
@@
The sheaves are locally free as we are working with locally trivial vector bundles and because the complex structure gives non-degenerate algebraic relations on the sections that smoothly depend on $M$. To prove the first claim, let $X\in\Gamma^{\omega}(M,E)$ be a chiral section and let $f\in\mathcal{O}^{\omega}_{M}$ be a chiral function. We compute
@@
If the spinorial representation $S^{1,0}$ underlying $\mathcal{F}^{1,0}$ carries an invariant non-degenerate graded symmetric bilinear form, the CS manifold $M^{1,0}$ carries a canonical volume form, denoted $\mathrm{vol}^{1,0}\in{\mathcal{B}}er_{M}^{1,0}$. The analogous statement holds for $M^{0,1}$.
@@ 5 Restriction to Ordinary Spacetimes
$i^{*}Q^{B}\times_{\check{M}}i^{*}Q^{F}\to\check{M}.$ (62)
The first factor defines the Riemannian structure, and the latter one defines the $\operatorname{Spin}^{\mathbb{F}}(V,q)$ structure. By definition of the superspace structure, there exists a map of principal bundles $Q^{F}\to Q^{B}$ with kernel $K$, and pulling back by $i^{*}$, we obtain the compatible structures
$i^{*}Q^{F}\to i^{*}Q^{B}\to\check{M}.$ (63)
Note that the pullback $i^{*}Q^{F}$, together with the spinorial representation $S$, allows for the construction of a spinor bundle over the ordinary spacetime
$S\check{M}=i^{*}Q^{F}\times_{\operatorname{Spin}^{\mathbb{F}}(V,q)}\Pi S\to\check{M}.$ (64)
@@
Finally, we have the following relations between the spinorial connection on $M$ and the spin connection on $i^{*}Q^{F}\to\check{M}$:
@@
Let $\varphi^{s}\in\mathcal{A}^{s}_{Q}$ be a spinorial connection. The pullback connection $i^{*}\varphi^{s}\in\mathcal{A}_{i^{*}Q}$ is a $\operatorname{Spin}^{\mathbb{F}}(V,q)$-connection on $i^{*}Q=\check{Q}$ (and thus in particular on the spinor bundle $S\check{M}$); and the pullback $i^{*}\left((\varphi^{s})^{\mathcal{B}}\right)\in\mathcal{A}_{i^{*}Q^{B}}$ of the $Q^{B}$-part of $\varphi^{s}$ is the Levi-Civita connection on $(\check{M},g|_{\check{M}})$.
@@ 6.1 The Construction
$\check{Q}\times T\Pi S\xrightarrow{\operatorname{pr}_{\Pi S}^{*}\mathrm{d}q}\tau^{*}\Pi S\check{M}\xhookrightarrow{}T\Pi S\check{M}.$ (84)
@@
$\delta_{\mathcal{F}}:\tau^{*}\Pi S\check{M}\to T\Pi S\check{M},\quad X\mapsto X+\frac{1}{2}\nabla^{\varphi}_{\Gamma(\mathcal{E},X)},$ (85)
@@
as after restriction, the structure sheaf is reduced to ordinary functions on $\check{M}$, where the covariant derivative reduces to the ordinary differential. The term involving $[\nabla_{\Gamma(\mathcal{E},\partial_{\mu})}^{\varphi},\nabla_{\Gamma(\mathcal{E},\partial_{\nu})}^{\varphi}]$ vanishes along the body as it has coefficients that are linear in $\Pi S$. Putting everything together, we find
@@
Let us fix (global) generators $\{\partial_{i}\}$ corresponding to an orthonormal basis $\{e_{i}\}$ of $V$, and $\{\partial_{\mu}\}$ corresponding to a basis $\{f_{\mu}\}$ of $\Pi S$. Let ${\Gamma_{\mu\nu}^{i}}e_{i}=\Gamma(f_{\mu},f_{\nu})$. In these coordinates, the structure distribution is given by
@@
such that $[D_{\mu},D_{\nu}]={\Gamma_{\mu\nu}}^{i}\partial_{i}$. This affine superspace constitutes a local model for general spinorial superspaces.
@@ 6.2 Computing Integrals
where the $(e_{i})$ are a local frame on $\check{M}$, i.e., a local section of $\check{Q}\to\check{M}$. In what follows, we identify frames and coframes by dualising the bases of the respective distributions. Note that there exists a $\operatorname{GL}(m|n)$ transformation of unit Berezinian relating $q$ with a frame that has the $\partial_{\mu}$ as odd generators (rather than $\delta_{\mathcal{F}}(\partial_{\mu}$), given explicitly by
@@
$\int_{M}f[q]=\int_{\check{M}}i^{*}(\partial_{1}\dots\partial_{n}f)e^{1}\wedge\dots\wedge e^{m}$ (105)
@@ Appendix B Frame Computations
Let $\psi$ be the sYM spinor associated to a connection $\mathsf{a}\in\mathcal{A}_{P}$, let $\varphi^{s}\in\mathcal{A}_{\hat{Q}}^{s}$ be a spinorial connection, and let $X$ be the horizontal lift of a section of the spinorial distribution $\mathcal{F}$ to $\mathcal{T}_{P}$. Set $\nabla^{\mathsf{a}}\eqqcolon\nabla$ and let $i:\check{M}\to M$ be the embedding of the ordinary spacetime manifold. It holds
@@
$\sqrt{n}\nabla_{X}\psi$ $=\iota_{X}\mathrm{d}^{\mathsf{a}}\langle\phi,\tilde{\gamma}\rangle=\langle\iota_{X}\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}\rangle,$ (143)
@@
$\sqrt{n}\nabla_{X}\psi$ $=\left\langle\iota_{X}\mathrm{d}^{\mathsf{a}}\phi,\sum_{Y}\tilde{\gamma}(Y^{\vee})\otimes Y\right\rangle$ (145)
$=\sum_{Y}\langle\iota_{Y}\iota_{X}\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}(Y^{\vee})\rangle$ (146)
@@
In the remainder of this appendix, we shall work out algebraic properties of component fields associated to a fixed connection $\mathsf{a}$ of the principal bundle $P\to M$ on the superspace. Therefore, we shall frequently write $F=F^{\mathsf{a}}$ for the curvature, as well as $f^{\mathsf{a}}=f$, $\psi^{\mathsf{a}}=\psi$ and $\phi^{\mathsf{a}}=\phi$ for the component fields.
@@
$=i^{*}\langle\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}\circ\gamma\circ\tilde{\epsilon}\rangle,$ (160)
@@
Let $\Delta_{\epsilon}^{\mathsf{a}}=\epsilon^{XY}\nabla_{X}\circ\nabla_{Y}$, let $a=i^{*}\mathsf{a}$, and let $\check{\psi}=i^{*}\psi$. It holds
$\big{(}\Delta^{\mathsf{a}}_{\epsilon}\psi\big{)}\big{|}_{\check{M}}=2D\!\!\!\!/\,^{a}\lambda.$ (161)
@@
$\sqrt{n}\cdot\Delta^{\mathsf{a}}_{\epsilon}\psi$ $=\epsilon^{XY}\iota_{X}\mathrm{d}^{\mathsf{a}}\iota_{Y}\mathrm{d}^{\mathsf{a}}\langle\phi,\tilde{\gamma}\rangle$ (162)
$=\epsilon^{XY}\langle\iota_{Z}\iota_{X}\mathrm{d}^{\mathsf{a}}\iota_{Y}\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}(Z^{\vee})\rangle$ (163)
$=\epsilon^{XY}\langle\iota_{X}\iota_{Z}L_{Y}\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}(Z^{\vee}),\rangle$ (164)
$=\epsilon^{XY}\langle\iota_{X}(\iota_{[Z,Y]}-\mathrm{d}^{\mathsf{a}}\iota_{Y}\iota_{Z})\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}(Z^{\vee})\rangle$ (165)
$=-\epsilon^{XY}\langle\iota_{[Z,Y]}\iota_{X}\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}(Z^{\vee})\rangle-\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{\mathsf{a}}\iota_{Y}\iota_{Z}\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}(Z^{\vee})\rangle$ (166)
@@
$\epsilon^{XY}\langle\iota_{[Z,Y]}\iota_{X}\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}(Z^{\vee})\rangle\big{|}_{\check{M}}$ $=\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}(Z^{\vee})\otimes\Gamma(Z,Y)\rangle\big{|}_{\check{M}}$ (167)
$=\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}\circ\gamma(Y)\rangle\big{|}_{\check{M}}$ (168)
$=-\langle\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}\circ\gamma\circ\tilde{\epsilon}\rangle\big{|}_{\check{M}}$ (169)
@@
$-\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{\mathsf{a}}\iota_{Y}\iota_{Z}\mathrm{d}^{\mathsf{a}}\phi,\tilde{\gamma}(Z^{\vee})\rangle\big{|}_{\check{M}}$ $=\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{\mathsf{a}}\iota_{[Y,Z]}f,\tilde{\gamma}(Z^{\vee})\rangle\big{|}_{\check{M}}$ (171)
@@
$\Delta^{\mathsf{a}}_{\epsilon}\psi\big{|}_{\check{M}}=2D\!\!\!\!/\,^{a}\lambda,$ (175)
@@
$\sqrt{n}\nabla_{X}\chi$ $=\iota_{X}\mathrm{d}^{\mathsf{a}}\langle\zeta,\tilde{\gamma}\rangle$ (179)
$=\sum_{Y}\langle\iota_{Y}\iota_{X}\mathrm{d}^{\mathsf{a}}\zeta,\tilde{\gamma}(Y^{\vee})\rangle$ (180)
$=-\sum_{Y}\langle\iota_{Y}\iota_{X}\mathrm{d}^{\mathsf{a}}(f+\bar{\zeta}),\tilde{\gamma}(Y^{\vee})\rangle$ (181)
$=-\sum_{Y}\langle\iota_{[Y,X]}\mathrm{d}^{\mathsf{a}}(f+\bar{\zeta}),\tilde{\gamma}(Y^{\vee})\rangle$ (182)
@@
$\sqrt{n}\nabla_{X}\chi=\iota_{X}\mathrm{d}^{\mathsf{a}}\langle\zeta,\tilde{\gamma}\rangle|_{\check{M}}=-\langle f+\bar{\zeta},\tilde{\gamma}\circ\gamma(X)\rangle|_{\check{M}}=-\langle f,\tilde{\gamma}\circ\gamma(X)\rangle|_{\check{M}}$ (183)
@@
$\big{(}\Delta^{\mathsf{a}}_{\epsilon}\chi\big{)}\big{|}_{\check{M}}\coloneqq\Big{(}\epsilon^{XY}\nabla_{X}\nabla_{Y}\chi\Big{)}\Big{|}_{\check{M}}=-D\!\!\!\!/\,^{a}\big{(}\chi|_{\check{M}}\big{)}.$ (185)
@@
$\sqrt{n}\cdot\Delta^{\mathsf{a}}_{\epsilon}\chi$ $=\epsilon^{XY}\iota_{X}\mathrm{d}^{\mathsf{a}}\iota_{Y}\mathrm{d}^{\mathsf{a}}\langle\zeta,\tilde{\gamma}\rangle$ (186)
$=\epsilon^{XY}\langle\iota_{Z}\iota_{X}\mathrm{d}^{\mathsf{a}}\iota_{Y}\mathrm{d}^{\mathsf{a}}\zeta,\tilde{\gamma}(Z^{\vee})\rangle$ (187)
$=\epsilon^{XY}\langle\iota_{X}\iota_{Z}L_{Y}\mathrm{d}^{\mathsf{a}}\zeta,\tilde{\gamma}(Z^{\vee}),\rangle$ (188)
$=\epsilon^{XY}\langle\iota_{X}(\iota_{[Z,Y]}-\mathrm{d}^{\mathsf{a}}\iota_{Y}\iota_{Z})\mathrm{d}^{\mathsf{a}}\zeta,\tilde{\gamma}(Z^{\vee})\rangle$ (189)
$=-\epsilon^{XY}\langle\iota_{[Z,Y]}\iota_{X}\mathrm{d}^{\mathsf{a}}\zeta,\tilde{\gamma}(Z^{\vee})\rangle-\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{\mathsf{a}}\iota_{Y}\iota_{Z}\mathrm{d}^{\mathsf{a}}\zeta,\tilde{\gamma}(Z^{\vee})\rangle$ (190)
@@
$\epsilon^{XY}\langle\iota_{[Z,Y]}\iota_{X}\mathrm{d}^{\mathsf{a}}\zeta,\tilde{\gamma}(Z^{\vee})\rangle\big{|}_{\check{M}}$ $=\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{\mathsf{a}}\zeta,\tilde{\gamma}(Z^{\vee})\otimes\Gamma(Z,Y)\rangle\big{|}_{\check{M}}$ (191)
$=\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{\mathsf{a}}\zeta,\tilde{\gamma}\circ\gamma(Y)\rangle\big{|}_{\check{M}}$ (192)
$=-\langle\mathrm{d}^{\mathsf{a}}\zeta,\tilde{\gamma}\circ\gamma\circ\tilde{\epsilon}\rangle\big{|}_{\check{M}}$ (193)
@@
$-\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{\mathsf{a}}\iota_{Y}\iota_{Z}\mathrm{d}^{\mathsf{a}}\zeta,\tilde{\gamma}(Z^{\vee})\rangle\big{|}_{\check{M}}$ $=\epsilon^{XY}\langle\iota_{X}\mathrm{d}^{\mathsf{a}}\iota_{[Y,Z]}(f+\bar{\zeta}),\tilde{\gamma}(Z^{\vee})\rangle\big{|}_{\check{M}}$ (195)